OpenStudy (lolacole12):

Which equation represents the word sentence? The product of 5 and a number increased by 7 is 32. A. 5x = 32 + 7 B. 5x + 32 = 7 C. 5x – 7 = 32 D. 5x + 7 = 32

OpenStudy (lolacole12):

fan and medal

OpenStudy (lolacole12):

@bestbruno

OpenStudy (lolacole12):

@lordhelix8th

OpenStudy (lolacole12):

@medjo

OpenStudy (lolacole12):

@Redneckk__Lifee

OpenStudy (lolacole12):

help

OpenStudy (anonymous):

D

OpenStudy (lolacole12):

thank you

OpenStudy (lolacole12):

Which is an equation for the given word problem? After Nancy deposits her birthday gift of $20.00 into her savings account, her balance is $225.50. Which equation can be used to find the amount, n, in her account before she made the deposit? A. 20.00 – n = 225.50 B. 20.00 + n = 225.50 C. n + 225.50 = 20.00 D. n – 225.50 = 20.00

OpenStudy (lolacole12):

@medjo

OpenStudy (anonymous):

ill do this one but can I get a fan and me

OpenStudy (lolacole12):

yes

OpenStudy (lolacole12):

@Redneckk__Lifee

OpenStudy (lolacole12):

@Gir_16

OpenStudy (lolacole12):

@♪Chibiterasu

OpenStudy (lolacole12):

:(

OpenStudy (anonymous):

go open your maths book kid !

OpenStudy (anonymous):

Its B NOW CAN

OpenStudy (lolacole12):

i already did @Redneckk__Lifee

OpenStudy (anonymous):

Okay thanks love!